OpenStudy (anonymous):

How do i simplify the following expression: 6(p - 6) + 9p

jimthompson5910 (jim_thompson5910):

Distribute to get 6(p - 6) + 9p 6p - 36 + 9p what's next?

OpenStudy (anonymous):

6(p-6)+9p = 6p - 36 +9p = 15p -36

OpenStudy (anonymous):

How did you get 36

jimthompson5910 (jim_thompson5910):

i multiplied the 6 outside the parenthesis by the -6 inside the parenthesis

OpenStudy (anonymous):

how do i do the same with : 0.1t+4.4v-2.7t

jimthompson5910 (jim_thompson5910):

no need to distribute, just need to combine like terms this time

jimthompson5910 (jim_thompson5910):

0.1t - 2.7t = ???

OpenStudy (anonymous):

i dont know a -2.6

jimthompson5910 (jim_thompson5910):

so -2.7t which means 0.1t+4.4v-2.7t simplifies to -2.6t+4.4v

OpenStudy (anonymous):

i need help with about 11 more problems i thought it was -2.6

jimthompson5910 (jim_thompson5910):

no there's a 't' at the end

OpenStudy (anonymous):

i know but you put -2.7t not -2.6t

jimthompson5910 (jim_thompson5910):

it starts out with -2.7t, but you combine it with 0.1t

jimthompson5910 (jim_thompson5910):

0.1t - 2.7t = -2.6t

OpenStudy (anonymous):

oo

jimthompson5910 (jim_thompson5910):

oh my bad, i know what you're saying...lol yeah i did, i meant to say -2.6t

OpenStudy (anonymous):

no prob so it is 2.6t +4.4v

jimthompson5910 (jim_thompson5910):

-2.6t + 4.4v

OpenStudy (anonymous):

ok next problem: 1/2(20m + 40)+ 1/2m

jimthompson5910 (jim_thompson5910):

distribute first, what do you get?

OpenStudy (anonymous):

ok what does distribute mean?

jimthompson5910 (jim_thompson5910):

take the 1/2 to the left of the parenthesis and multiply it with everything inside the parenthesis

jimthompson5910 (jim_thompson5910):

1/2 times 20m = 10m 1/2 times 40 = 20

jimthompson5910 (jim_thompson5910):

so after distributing, you get 10m + 20 + 1/2m

OpenStudy (anonymous):

so it would be .5 * 40

jimthompson5910 (jim_thompson5910):

you can write it like that (since 1/2 = 0.5)

OpenStudy (anonymous):

do i add the two products

jimthompson5910 (jim_thompson5910):

only if they are like terms

OpenStudy (anonymous):

i got: 10.5m + 20

jimthompson5910 (jim_thompson5910):

good, you got it if you want to convert that to fraction form (you don't have to since your answer is just fine), you'll get \[\Large \frac{21}{2}m+20\]

OpenStudy (anonymous):

next problem: 13/4x+ 1/8 - 1/2x+ 3/4

jimthompson5910 (jim_thompson5910):

combine 13/4x and -1/2x, what do you get?

jimthompson5910 (jim_thompson5910):

I would keep things as fractions, but you can convert to decimals if you want

OpenStudy (anonymous):

that is a subtraction sign before 1/2x

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

so 13/4x minus 1/2x

OpenStudy (anonymous):

but there is a addition sign after 13/4x

jimthompson5910 (jim_thompson5910):

well you're not adding 1/8 to 13/4x though (since they aren't like terms)

OpenStudy (anonymous):

i got : 30/8

jimthompson5910 (jim_thompson5910):

\[\Large \frac{13}{4}x-\frac{1}{2}x\] \[\Large \frac{13}{4}x-\frac{2}{4}x\] \[\Large \frac{13-2}{4}x\] \[\Large \frac{11}{4}x\]

jimthompson5910 (jim_thompson5910):

So \[\Large \frac{13}{4}x-\frac{1}{2}x\] combines to \[\Large \frac{11}{4}x\]

OpenStudy (anonymous):

oo i added

jimthompson5910 (jim_thompson5910):

i see, stick with the sign that's directly in front of the 1/2 (which is a - sign)

OpenStudy (anonymous):

ok wats the next step

jimthompson5910 (jim_thompson5910):

combine 1/8 and 3/4

OpenStudy (anonymous):

after this i only have one more question for understanding

OpenStudy (anonymous):

i got 7/8

OpenStudy (anonymous):

wat do i do with the multiplication sign

jimthompson5910 (jim_thompson5910):

good, so \[\Large \frac{13}{4}x+\frac{1}{8}-\frac{1}{2}x+\frac{3}{4}\] simplifies to \[\Large \frac{11}{4}x+\frac{7}{8}\]

OpenStudy (anonymous):

never mind i got confused about the multiplication sign\

OpenStudy (anonymous):

thanxs for all

jimthompson5910 (jim_thompson5910):

yw